OpenStudy (anonymous):

Johnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (lgbasallote):

let b = boys g = girls b + g = 82 b = 16 + 2g can you do it from there?

OpenStudy (anonymous):

16+2g+g=82?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

3g=66?

OpenStudy (anonymous):

g=22?!?!